Solve :
(D²-2DD´+D´²)z = 12xy
1
Expert's answer
2019-08-28T05:02:16-0400

Consider a quadratic form "x^2-2xy+y^2" (for a differential equation "A_{11}\frac{\partial^2 z}{\partial x^2}+A_{12}\frac{\partial^2 z}{\partial x\partial y}+A_{22}\frac{\partial^2 z}{\partial y^2}+\ldots=0" we consider a quadratic form "A_{11}x^2+A_{12}xy+A_{22}y^2" ). Its matrix is "\begin{pmatrix}
 1 & -1 \\
 -1 & 1
\end{pmatrix}" with eigenvectors "(1,1)" and "(1,-1)" .

So we have change of the variables: "u=1\cdot x+1\cdot y = x+y" and "v=1\cdot x+(-1)\cdot y = x-y" .

Then "\frac{\partial f}{\partial x}=\frac{\partial f}{\partial u}\frac{\partial u}{\partial x}+\frac{\partial f}{\partial v}\frac{\partial v}{\partial x}=\frac{\partial f}{\partial u}+\frac{\partial f}{\partial v}" and "\frac{\partial f}{\partial y}=\frac{\partial f}{\partial u}\frac{\partial u}{\partial y}+\frac{\partial f}{\partial v}\frac{\partial v}{\partial y}=\frac{\partial f}{\partial u}-\frac{\partial f}{\partial v}" . Move from the variables "x,y" to the variables "u,v" .

In case "f=z" we have "\frac{\partial z}{\partial x}=\frac{\partial z}{\partial u}+\frac{\partial z}{\partial v}" and "\frac{\partial z}{\partial y}=\frac{\partial z}{\partial u}-\frac{\partial z}{\partial v}" .

In case "f=\frac{\partial z}{\partial x}" we have "\frac{\partial^2 z}{\partial x^2}=\frac{\partial}{\partial x}\bigl(\frac{\partial z}{\partial x}\bigr)=\frac{\partial}{\partial u}\bigl(\frac{\partial z}{\partial x}\bigr)+\frac{\partial}{\partial v}\bigl(\frac{\partial z}{\partial x}\bigr)=\frac{\partial^2 z}{\partial u^2}+2\frac{\partial^2 z}{\partial u\partial v}+\frac{\partial^2 z}{\partial v^2}" and "\frac{\partial^2 z}{\partial x\partial y}=\frac{\partial}{\partial y}\bigl(\frac{\partial z}{\partial x}\bigr)=\frac{\partial}{\partial u}\bigl(\frac{\partial z}{\partial x}\bigr)-\frac{\partial}{\partial v}\bigl(\frac{\partial z}{\partial x}\bigr)=\frac{\partial^2 z}{\partial u^2}-\frac{\partial^2 z}{\partial v^2}"

In case "f=\frac{\partial z}{\partial y}" we have "\frac{\partial^2 z}{\partial y^2}=\frac{\partial}{\partial y}\bigl(\frac{\partial z}{\partial y}\bigr)=\frac{\partial}{\partial u}\bigl(\frac{\partial z}{\partial y}\bigr)-\frac{\partial}{\partial v}\bigl(\frac{\partial z}{\partial y}\bigr)=\frac{\partial^2 z}{\partial u^2}-2\frac{\partial^2 z}{\partial u\partial v}+\frac{\partial^2 z}{\partial v^2}" .

In addition "x=\frac{u+v}{2}" and "y=\frac{u-v}{2}" .

Rewrite the original equation: "\bigl(\frac{\partial^2 z}{\partial u^2}+2\frac{\partial^2 z}{\partial u\partial v}+\frac{\partial^2 z}{\partial v^2}\bigr)-2\bigl(\frac{\partial^2 z}{\partial u^2}-\frac{\partial^2 z}{\partial v^2}\bigr)+"

"+\bigl(\frac{\partial^2 z}{\partial u^2}-2\frac{\partial^2 z}{\partial u\partial v}+\frac{\partial^2 z}{\partial v^2}\bigr)=12\frac{u+v}{2}\frac{u-v}{2}" . That is "\frac{\partial^2 z}{\partial v^2}=\frac{3}{4}u^2-\frac{3}{4}v^2"

Integrate it with respect to "v" : "\frac{\partial z}{\partial v}=\frac{3}{4}u^2v-\frac{1}{4}v^3+R(u)" and again "z=\frac{3}{8}u^2v^2-\frac{1}{16}v^4+R(u)v+S(u)",

where "R(u), S(u)" are arbitrary twice differentiable functions.

Return to the variables "x" and "y" : "z=\frac{3}{8}(x+y)^2(x-y)^2-\frac{1}{16}(x-y)^4+(x-y)\cdot R(x+y)+"

"+S(x+y)"